OpenStudy (anonymous):

a line has slope -2 and contains P(3,4) and Q(-4,a). find the value of a and show work :p

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ P&({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}})\quad Q&({\color{red}{ -4}}\quad ,&{\color{blue}{ a}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}=-2\) solve for "a"

OpenStudy (anonymous):

i got a-4 all over -7? what did u get @jdoe0001

OpenStudy (jdoe0001):

that's correct.. but you'd still need to solve for "a" so cross-multiply and add 4 to both sides

OpenStudy (anonymous):

?what do u mean by that

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ P&({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}})\quad Q&({\color{red}{ -4}}\quad ,&{\color{blue}{ a}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ a}}-{\color{blue}{ 4}}}{{\color{red}{ -4}}-{\color{red}{ 3}}}=-2\implies \cfrac{a-4}{{\color{red}{ -7}}}=-2\\ \quad \\ a-4={\color{red}{ -7}}\cdot 2\implies a\cancel{-4+4}=(7\cdot 2)+4\)

OpenStudy (anonymous):

so a=18?

OpenStudy (jdoe0001):

so a=18? \(\large \checkmark\)

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

yw